Find the ending balance in an account that opens with $3,590, earns 5.5% interest compounded monthly, and is held for 15 years. (round answer to nearestcent)
Answer by cta1987(8)

A=P(1+r/n)^nt
monthly=12 times per year
rate=0.055
P=3590
t=15 years
A=3590(1+0.055/12)^12*15
A=3590(1+0.0458333333)^180
A=3590(1.0458333333)^180
A=3590(3186.0041424)
A=11437754.79